OpenStudy (anonymous):

find all solutions to the equation 7 sin^2 x - 14 sin x + 2 = -5

OpenStudy (anonymous):

call sin x = x. Solve the quadratic equation: 7x^2 - 14x + 7 = 0 x^2 - 2x + 1 = 0 since a + b + c = 0, one real root is 1 and the other is 1/1 = 1. sin x = 1 = sin Pi/2 --> x = Pi/2